OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. x2 - 4x - 32 = 0

OpenStudy (anonymous):

x=-4 x=8

OpenStudy (anonymous):

http://www.youtube.com/watch?v=xGOQYTo9AKY x^2-4x-32 = 0 (x^2-4x+4) -32 -4 = 0 (x-2)(x-2) -36 = 0 (x-2)^2 = 36 sqrt((x-2)^2) = sqrt(36) x-2 = +/-6 Where x-2 = +6 x = 8 Where x-2 = -6 x=-4